OpenStudy (anonymous):

Need help with differential equations :/. Find a second solution for xy''+(x-1)y'-y=0, given that y1=e^-x.

OpenStudy (anonymous):

Do you know how to use integrating factor?

OpenStudy (anonymous):

Oh... yes I do but I thought that only worked if the equation was in standard form and a first order equation. y'+P(x)y=q(x)

OpenStudy (anonymous):

In that case I would just multiply everything by e^(integral (p) dx). Does that work on this too?

OpenStudy (anonymous):

Ah wait do I need to do reduction of order to get it in the right form?

OpenStudy (anonymous):

Oh wait, give me a few minutes. I'm not sure if it applies here.

OpenStudy (anonymous):

okay thank you. I appreciate it because I am definitely lost :/.

OpenStudy (anonymous):

I actually am taking Differential Equations right now for my engineering major, but we haven't covered second order differential equations with nonconstant coefficients. I would recommend however visiting this page http://tutorial.math.lamar.edu/Classes/DE/ReductionofOrder.aspx It seems to be close to what you are looking for. :)

OpenStudy (anonymous):

Okay I will read over it right now. Thank you for the link :). and that is interesting. I studied molecular biology but I am switching to biosystems engineering.

OpenStudy (anonymous):

If they were constants, I'd totally be able to help you using the different cases :) I can't do anything biology for the life of me, haha. I admit I find it extremely fascinating but I can never memorize all the terms and functions XD

OpenStudy (anonymous):

hmm... okay so I did find an answer key to the question but I am still not sure how the professor solved this. Apparently the answer is y2=x-1 but I am still trying to figure out what she did.

OpenStudy (anonymous):

I think I understand the basic idea a solution to the differential equation can be a scalar multiple of the given y1. but I don't understand where the integrating factor comes in... this is so frustrating :/.

OpenStudy (anonymous):

I don't understand why he used integrating factor at first. It seems like a different method than reduction of order, but I actually understand what he's doing. It seems like after finding e^-u, he integrating in the following form \[y_2=y_1* \int\limits \frac{ e^{-\mu} }{ y_1{^2} }dx\] Then once he simplified that, he just plugged in y1 that was given from the beginning. Also note that at the end, he put an e^-x instead of e^x for the last term in the parentheses. Or vice versa. Not sure which one it's suppose to be but there's an error there.

OpenStudy (anonymous):

Does this make sense now?

OpenStudy (anonymous):

I sort of understand but I don't really get why the professor is multiplying by ∫e^(−μ)/y^2)dx. Where does that come from?

OpenStudy (anonymous):

I'm assuming it's some sort of formula to follow. I wouldn't really know because I haven't learned this yet (until now though, so thank you :P). But I don't understand the logic behind it and where it was derived from. For now I would just go with it and ask your professor later for clarification.

OpenStudy (anonymous):

haha... okay... well, if it works on the exam, I won't question it lol.

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Thanks for the help :). Good luck with your studies as well. Hopefully we can help eachother out :p

OpenStudy (anonymous):

You are very welcome! And hopefully :P